OpenStudy (anonymous):

This limit represents the derivative of a function f(x) at a particular number in the domain: lim as X->4 (x^3-64)/(x-4) a) Find f(x) b) Find a Answer: 4 c) Find the slope of the tangent line to the graph of f(x) at x=a

OpenStudy (anonymous):

function is the cubing function \(f(x)=x^3\)

OpenStudy (anonymous):

Can you explain how you got that? and that means the slope would be? \[f \prime (x)=3x^2 \] so \[f \prime (4) = 3*4^2\] \[=48\]